# Emergent Coordination in Multi-Agent Language Models

### v2

@@ 2 Method
To explore system dynamics, we implement a coalition test, an extension of the practical criterion to triplets. Let $I_{3}$ be the mutual information between a triplet’s current state and the future macro signal: $I_{3}=I((X_{i,t},X_{j,t},X_{k,t});V_{t+\ell})$. I.e., how much the three agents jointly predict the macro’ future serves as a measure of the system’s information-processing capacity (Mediano et al., 2022a). High $I_{3}$ indicates that agents’ behaviors are tightly coupled and contain rich information about the emerging collective outcome, a pattern frequently associated with collective intelligence (De Vincenzo et al., 2017; Wicks et al., 2007; Mora and Bialek, 2011; Barnett et al., 2013). Low $I_{3}$ signals weak alignment, resulting in uncoordinated or chaotic collective behavior. Then
@@
We use the Williams–Beer two‑source PID (Williams and Beer, 2010) with the $I_{min}$ (minimum specific information) redundancy, estimated via plug‑in probabilities. For robustness, we also use Jeffreys-smoothed probabilities, and alternatively MMI redundancy (minimum of whole mutual informations; Mediano et al., 2025) with mutual information estimated with Miller–Madow bias correction (Miller, 1955). All variables are discretized into $K=2$ bins using quantile binning. Data are encoded as factors with fixed levels 1,2{1,2} (and for the joint target as the Cartesian product levels) to avoid dropping empty categories. We report $\ell=1$ which is suitable to detect next-step oscillating behavior (see task description). Because trials end at success, episode lengths vary (see plot of rounds to success in Appendix). We report additional sensitivity tests using “early synergy” truncated to a fixed horizon of $H\in\{10,15\}$ rounds to ensure comparability (results in Appendix A.2 and A.12; main analyses use full trajectories).
@@
Small-sample entropy estimation is often challenging because the true distribution is only partially observed and many outcome categories receive few or even zero counts (“empty bins”). Such finite-sample estimation of information measures is biased upward and can yield spurious positive findings. Bias grows with more bins, dimensionality (e.g., using triplets instead of pairs), and small $N$ (few timesteps). The empirical plug-in estimator replace true probabilities with sample frequencies and unseen (or under-sampled) outcomes often are assigned zero or very low probabilities (Hausser and Strimmer, 2009). To account for this issue, we follow best practices and take several steps. First, we compute the emergence capacity and the coalition tests as order $k=2$ instead of the entire system of $n$ agents. Second, we apply quantile binning (with two bins) to reduce the dimensionality of our data (sensitivity analyses with three bins reported in A.8). Third, we use bias-corrected entropy estimator with Jeffreys’ prior, which smooths estimates and avoids empty bins by adding $\alpha=\frac{1}{2}$ pseudo-counts in the Dirichlet estimator (Jeffreys, 1946). This reduces systematic inflation and cures zero‑count pathologies. Finally, as a robustness check, we also report results using the Miller–Madow bias-corrected estimator (Miller, 1955), and the MMI redundancy measure (MMI typically overestimates redundancy, making synergy estimates more conservative; Mediano et al., 2025).
@@ 4.2 Dynamical Mechanisms
Decomposition this stability reveals the specific mechanism of coordination. While the system exhibits emergence capacity and practical emergence, we find the stability in group behavior is supported by pairwise alignment rather than irreducible triplet complexity. Specifically, triadic information gain ($G_{3}$) is around 0 in Persona and ToM (Figure 3c). The Plain conditions shows small positive $G_{3}>0$ (Wilcoxon $p$-value = $0.026$) but given the near-zero Total Stability, this likely reflects transient stochastic correlations and oscillation rather than sustained coordination. In dynamical systems terms, the ToM prompt creates a deep basin of attraction where agents converge on a shared schematic (or Schelling point), making the collective state robust but informationally redundant. This mirrors “Mean Field” coupling in physics: because agents receive only global feedback and cannot observe individual peers, they couple to the aggregate signal rather than forming distinct local bonds. In such an environment, attempting complex higher-order synergy would be fragile; instead, the system converges on the efficient solution of dense pairwise alignment to the Mean Field.
@@
Next, we explore whether agents evolve specialized roles and identities. Are there detectable between-agent differences in either the level of their contribution to the group guess or their temporal evolution (i.e., learning rate)? Using the hierarchical model comparison test we find that agents in may groups have differentiated identities (Figure 3a; a Fisher test for joint $p$-value evidence is again highly significant). There is substantially more differentiation among agents in the Persona condition, and even more in the ToM condition. Inspecting the model output of agents, they often contain references to the “personal experience” of their assigned persona such as “In my experience (whether it’s corraling cattle or wrangling numbers), starting toward the middle gives the group the best shot to zero in after the first feedback” (see Appendix for additional sample quotes). In Appendix A.10 we provide additional robustness tests ruling out heterogeneous learning rates as the sole source of differentiation and provide additional evidence for stable agent identities that persist over time. That is, in the Plain condition, differentiation results only from idiosyncratic noise and transient drift introduced by probabilistic LLM responses. With Persona, agents gain stable, identity linked behavioral preferences that refine their differentiation. The addition, ToM sharpens role differentiation through feedback loops from mutual predictive modeling by conditioning on the public history (which functions as a common ground and coordination device; Lewis, 1969; Skyrms, 2010). This converts small persona-induced asymmetries into stable, self-reinforcing roles, creating the basis for strategic complementarity and stable alignment (answering RQ2 and RQ3)
@@ A.11 Additional Details for Performance Analysis
We address these issue as follows. First, we focus our analysis on the interaction of synergy and redundancy, rather than each variable individually to capture the tradeoff between the two. Second, we compute both emergent properties (synergy and redundancy) only on early rounds (in our case, 10 time steps). We chose 10 rounds as a good tradeoff between giving the multi-agent systems enough time to evolve, while limiting the amount of runs that have already completed. While this reduces our ability to detect the full emergent potential of the multi-agent system (since agents have less time to evolve) it creates a clean apples-to-apples comparison between measured values (they are all based on exactly 10 time steps). In that sense, our analyses likely underestimate the full effect of synergy because we only capture the early onset of synergy. Focusing on only the first 10 time steps, leads to additional complications, however. About 23% of experiments finish early and do not make it till round ten (roughly balanced across treatments with 22% in Plain, 25% in Persona, and 20% in ToM). Simply removing these observations would bias our results since this censoring occurs post-treatment assignment: removing those observations would introduce sampling bias. Instead, we take these observations as censored (i.e., we cannot compute the early synergy measure) and control for this censoring using stabilized inverse probability weighting (IPW; Wooldridge, 2002). That is, in a separate modeling step we predict (via logistic regression and only task difficulty as exogenous predictor variables that is independent of the treatment) the likelihood of observing (early) synergy and redundancy. [^5] As measures for task difficulty we use (a) how far groups target number was from the mid point (smaller numbers requiring more steps of the binary search process); and (b) the modulo remainder of the target number and the group size (target numbers that are closer to divisibility by the group size are easier to guess by pure chance). We then take the inverse of the predicted likelihood as weight in the causal mediation analysis. This adjust observations in the primary analysis with the inverse of the probability of making such observations.
@@ A.13 Experiments with other Models
As a reasoning model, Qwen3 exhibited persistent looping behavior during its chain-of-thought process—a known failure mode of reasoning models (Pipis et al., 2025). These models are particularly susceptible to self-reflection and deductive reasoning traps (Baek and Tegmark, 2025). In our setting, we identify three specific patterns of oscillatory reasoning that arise reliably as soon as group feedback becomes inconsistent with the individual binary search strategy.
@@
These examples illustrate that the model is aware of the task ambiguity and often correctly recognizing the oscillating pattern (often, reasoning trace includes the word phase “[we are] oscillating”). That is, the failure is not due to reasoning error. Instead it is failure to terminate under unresolved ambiguity. We term this paralysis under coordination ambiguity. Notably, the paralysis persists even at a high temperature setting ($T=1.0$), which should increase stochastic exploration and disrupt deterministic decoding loops, indicating that the instability arises from internal reasoning dynamics rather than sampling artifacts. As a practical remedy, we managed to improve robustness by adding one extra line to the prompt to to the Qwen experiments:
@@ A.14 Specific Persona Effects
Both the Persona and ToM conditions rely on personas. Do the specific personas—or combinations of personas—present in a multi-agent system affect performance or the emergence of higher-order structures? We explore these questions with several test. Recall that our experiments are based on 20 different personas that are generated once globally, and then reused throughout the experiments by sampling 10 persona from the library (without replacement). First, we perform joint F-tests (based on linear regression) of the null hypothesis that all persona dummies have no effect on the outcome (i.e., no specific persona substantially affects group outcomes and all coefficients equal zero). Rejecting the null would indicate that at least one of the persona indicators is significantly associated with the outcome variable. We find no evidence that the presence of specific personas effect performance ($F=1.217$; $p=0.239$), number of rounds to success ($F=1.356$; $p=0.144$), PID synergy ($F=0.839$; $p=0.66$), or PID redundancy ($F=0.596$; $p=0.910$). These results suggest that what matters is the presence of personas overall—markers that allow the LLM agents to center their differentiation on—rather than who those personas are specifically.

### v3

@@ 2 Method
To explore system dynamics, we implement a coalition test, an extension of the practical criterion to triplets. Let $I_{3}$ be the mutual information between a triplet’s current state and the future macro signal: $I_{3}=I((X_{i,t},X_{j,t},X_{k,t});V_{t+\ell})$. I.e., how much the three agents jointly predict the macro’s future serves as a measure of the system’s information-processing capacity (Mediano et al., 2022a). High $I_{3}$ indicates that agents’ behaviors are tightly coupled and contain rich information about the emerging collective outcome, a pattern frequently associated with collective intelligence (De Vincenzo et al., 2017; Wicks et al., 2007; Mora and Bialek, 2011; Barnett et al., 2013). Low $I_{3}$ signals weak alignment, resulting in uncoordinated or chaotic collective behavior. Then
@@
We use the Williams–Beer two‑source PID (Williams and Beer, 2010) with the $I_{min}$ (minimum specific information) redundancy, estimated via plug-in probabilities. For robustness, we also use Jeffreys-smoothed probabilities, and alternatively MMI redundancy (minimum of whole mutual informations; Mediano et al., 2025) with mutual information estimated with Miller–Madow bias correction (Miller, 1955). All variables are discretized into $K=2$ bins using quantile binning. Data are encoded as factors with fixed levels 1,2{1,2} (and for the joint target as the Cartesian product levels) to avoid dropping empty categories. We report $\ell=1$ which is suitable to detect next-step oscillating behavior (see task description). Because trials end at success, episode lengths vary (see plot of rounds to success in Appendix). We report additional sensitivity tests using “early synergy” truncated to a fixed horizon of $H\in\{10,15\}$ rounds to ensure comparability (results in Appendix A.2 and A.12; main analyses use full trajectories).
@@
Small-sample entropy estimation is often challenging because the true distribution is only partially observed and many outcome categories receive few or even zero counts (“empty bins”). Such finite-sample estimation of information measures is biased upward and can yield spurious positive findings. Bias grows with more bins, dimensionality (e.g., using triplets instead of pairs), and small $N$ (few timesteps). The empirical plug-in estimator replaces true probabilities with sample frequencies and unseen (or under-sampled) outcomes often are assigned zero or very low probabilities (Hausser and Strimmer, 2009). To account for this issue, we follow best practices and take several steps. First, we compute the emergence capacity and the coalition tests as order $k=2$ instead of the entire system of $n$ agents. Second, we apply quantile binning (with two bins) to reduce the dimensionality of our data (sensitivity analyses with three bins reported in A.8). Third, we use bias-corrected entropy estimator with Jeffreys’ prior, which smooths estimates and avoids empty bins by adding $\alpha=\frac{1}{2}$ pseudo-counts in the Dirichlet estimator (Jeffreys, 1946). This reduces systematic inflation and cures zero‑count pathologies. Finally, as a robustness check, we also report results using the Miller–Madow bias-corrected estimator (Miller, 1955), and the MMI redundancy measure (MMI typically overestimates redundancy, making synergy estimates more conservative; Mediano et al., 2025).
@@ 4.2 Dynamical Mechanisms
Decomposing this stability reveals the specific mechanism of coordination. While the system exhibits emergence capacity and practical emergence, we find the stability in group behavior is supported by pairwise alignment rather than irreducible triplet complexity. Specifically, triadic information gain ($G_{3}$) is around 0 in Persona and ToM (Figure 3c). The Plain conditions shows small positive $G_{3}>0$ (Wilcoxon $p$-value = $0.026$) but given the near-zero Total Stability, this likely reflects transient stochastic correlations and oscillation rather than sustained coordination. In dynamical systems terms, the ToM prompt creates a deep basin of attraction where agents converge on a shared schematic (or Schelling point), making the collective state robust but informationally redundant. This mirrors “Mean Field” coupling in physics: because agents receive only global feedback and cannot observe individual peers, they couple to the aggregate signal rather than forming distinct local bonds. In such an environment, attempting complex higher-order synergy would be fragile; instead, the system converges on the efficient solution of dense pairwise alignment to the Mean Field.
@@
Next, we explore whether agents evolve specialized roles and identities. Are there detectable between-agent differences in either the level of their contribution to the group guess or their temporal evolution (i.e., learning rate)? Using the hierarchical model comparison test we find that agents in many groups have differentiated identities (Figure 3a; a Fisher test for joint $p$-value evidence is again highly significant). There is substantially more differentiation among agents in the Persona condition, and even more in the ToM condition. The reasoning traces of agents often contain references to the “personal experience” of their assigned persona such as “In my experience (whether it’s corraling cattle or wrangling numbers), starting toward the middle gives the group the best shot to zero in after the first feedback” (see Appendix for additional sample quotes). In Appendix A.10 we provide additional robustness tests ruling out heterogeneous learning rates as the sole source of differentiation and provide additional evidence for stable agent identities that persist over time. That is, in the Plain condition, differentiation results only from idiosyncratic noise and transient drift introduced by probabilistic LLM responses. With Persona, agents gain stable, identity linked behavioral preferences that refine their differentiation. The addition, ToM sharpens role differentiation through feedback loops from mutual predictive modeling by conditioning on the public history (which functions as a common ground and coordination device; Lewis, 1969; Skyrms, 2010). This converts small persona-induced asymmetries into stable, self-reinforcing roles, creating the basis for strategic complementarity and stable alignment (answering RQ2 and RQ3)
@@ A.11 Additional Details for Performance Analysis
We address these issues as follows. First, we focus our analysis on the interaction of synergy and redundancy, rather than each variable individually to capture the tradeoff between the two. Second, we compute both emergent properties (synergy and redundancy) only on early rounds (in our case, 10 time steps). We chose 10 rounds as a good tradeoff between giving the multi-agent systems enough time to evolve, while limiting the amount of runs that have already completed. While this reduces our ability to detect the full emergent potential of the multi-agent system (since agents have less time to evolve) it creates a clean apples-to-apples comparison between measured values (they are all based on exactly 10 time steps). In that sense, our analyses likely underestimate the full effect of synergy because we only capture the early onset of synergy. Focusing on only the first 10 time steps, leads to additional complications, however. About 23% of experiments finish early and do not make it till round ten (roughly balanced across treatments with 22% in Plain, 25% in Persona, and 20% in ToM). Simply removing these observations would bias our results since this censoring occurs post-treatment assignment: removing those observations would introduce sampling bias. Instead, we take these observations as censored (i.e., we cannot compute the early synergy measure) and control for this censoring using stabilized inverse probability weighting (IPW; Wooldridge, 2002). That is, in a separate modeling step we predict (via logistic regression and only task difficulty as exogenous predictor variables that is independent of the treatment) the likelihood of observing (early) synergy and redundancy. [^5] As measures for task difficulty we use (a) how far groups target number was from the mid point (smaller numbers requiring more steps of the binary search process); and (b) the modulo remainder of the target number and the group size (target numbers that are closer to divisibility by the group size are easier to guess by pure chance). We then take the inverse of the predicted likelihood as weight in the causal mediation analysis. This adjust observations in the primary analysis with the inverse of the probability of making such observations.
@@ A.13 Experiments with other Models
As a reasoning model, Qwen3 exhibited persistent looping behavior during its chain-of-thought process—a known failure mode of reasoning models (Pipis et al., 2025; Cemri et al., 2025; Shapira et al., 2026). These models are particularly susceptible to self-reflection and deductive reasoning traps (Baek and Tegmark, 2025) and circular reasoning (Duan et al., 2026). In our setting, we identify three specific patterns of oscillatory reasoning that arise reliably as soon as group feedback becomes inconsistent with the individual binary search strategy.
@@
These examples illustrate that the model is aware of the task ambiguity and often correctly recognizing the oscillating pattern (often, reasoning trace includes the word phase “[we are] oscillating”). That is, the failure is not due to reasoning error. Instead it is failure to terminate under irreducible epistemic uncertainty about others. We term this paralysis under coordination ambiguity. Notably, the paralysis persists even at a high temperature setting ($T=1.0$), which should increase stochastic exploration and disrupt deterministic decoding loops, indicating that the instability arises from internal reasoning dynamics rather than sampling artifacts. As a practical remedy, we managed to improve robustness by adding one extra line to the prompt to to the Qwen experiments:
@@ A.14 Specific Persona Effects
Both the Persona and ToM conditions rely on personas. Do the specific personas—or combinations of personas—present in a multi-agent system affect performance or the emergence of higher-order structures? We explore these questions with several test. Recall that our experiments are based on 20 different personas that are generated once globally, and then reused throughout the experiments by sampling 10 personas from the library (without replacement). First, we perform joint F-tests (based on linear regression) of the null hypothesis that all persona dummies have no effect on the outcome (i.e., no specific persona substantially affects group outcomes and all coefficients equal zero). Rejecting the null would indicate that at least one of the persona indicators is significantly associated with the outcome variable. We find no evidence that the presence of specific personas effect performance ($F=1.217$; $p=0.239$), number of rounds to success ($F=1.356$; $p=0.144$), PID synergy ($F=0.839$; $p=0.66$), or PID redundancy ($F=0.596$; $p=0.910$). These results suggest that what matters is the presence of personas overall—markers that allow the LLM agents to center their differentiation on—rather than who those personas are specifically.